# Lin Xu, Xin Zhong

### v1

@@ 1 Introduction
First, motivated by [13], we attempt to obtain the useful dissipation estimates from the basic energy inequality in terms of $L_{t}^{\infty}L_{x}^{3}$-norm of the density and $L_{t}^{2}L_{x}^{2}$-norm of $\nabla\theta$ (see (3.1)). With the help of a priori hypothesis (3.2) and the structure of system (1.1), we derive an estimate on the $L_{t}^{\infty}L_{x}^{3}$-norm of density (see Lemma 3.2). The next step is to obtain an estimate on the $L_{t}^{2}L_{x}^{2}$-norm of the gradient of temperature. At this stage, our argument relies on the temperature equation instead of the total energy equation used in [17]. This requires us to control the quantities $\sqrt{\rho}\dot{u}$ and $\nabla^{2}b$ in $L_{t}^{2}L_{x}^{2}$ and $\nabla b$ in $L_{t}^{\infty}L_{x}^{2}$ (see Lemma 3.3). Multiplying (1.1)2 by $u_{t}$ and applying delicate energy estimates, the estimate on the $L_{t}^{2}L_{x}^{2}$-norm for $\sqrt{\rho}\dot{u}$ can be obtained. It should be emphasized that, due to the term $\operatorname{curl}b\times b$, the estimate for $\sqrt{\rho}\dot{u}$ generates the complicated quantity $K_{2}(t)$ (see (3.4)). To control all these magnetic effects, one needs appropriate estimates for the magnetic field, which are derived from (1.1)4 (see Lemma 3.5).
Next, the key issue is to enclose the bootstrap argument. Based on the estimates derived in Lemmas 3.13.5, one obtains the uniform bound of density (via characteristic arguments) and the desired control of the quantity $S_{t}$ provided that the initial scaling-invariant quantity $N_{0}$ is sufficiently small (see Corollaries 3.1 and 3.2). In particular, in order to construct a complete scaling-invariant quantity that meets the requirements of the a priori estimates and the bootstrap argument, we multiply (3) by $(1+\bar{\rho}+\tfrac{1}{\bar{\rho}})$. Moreover, we succeed in establishing the estimate for $\|\rho\|_{L^{3}}$ in (3). During this process, because of the strong nonlinear coupling between the fluid velocity, temperature, and magnetic field, the key techniques in [23] cannot be applied in a straightforward manner. A new ingredient in our analysis is the introduction of the quantity
$(\bar{\rho}^{2}+\bar{\rho})\sup_{t\in[0,T]}\|\rho(t)\|_{L^{3}}\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}\,dt,$
which is assumed to be sufficiently small (see (3.2)). This structure allows us to close the estimate of $\|\rho\|_{L^{3}}$ and completes the bootstrap argument when the condition (1.8) holds.
@@ 4. $(d)$
Under the hypotheses of Theorem 1.1, $\epsilon_{0}>0$ is the small positive constant that appears in (1.8). Then there exist positive constants $C_{1}$ and $\epsilon_{2}$, depending only on the physical parameters (but independent of the initial data, $\bar{\rho}$, and $T$), with $\epsilon_{2}$ further being determined by (3.14), (3.21), (3.23), (3.34), (3.39), (3.44), (3), (3), and (3), such that if for all $t\in[0,T]$,
$\rho\leq 2\bar{\rho},\quad S_{t}+(\bar{\rho}^{2}+\bar{\rho})\sup_{t\in[0,T]}\|\rho(t)\|_{L^{3}}\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}\,dt\leq 2\epsilon,$ (3.2)
@@
$\rho\leq\frac{3}{2}\bar{\rho},\quad S_{t}+(\bar{\rho}^{2}+\bar{\rho})\sup_{t\in[0,T]}\|\rho(t)\|_{L^{3}}\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}\,dt\leq\frac{3}{2}\epsilon.$ (3.3)
@@
$\|\sqrt{\rho}\theta(t)\|_{L^{2}}^{2}+\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}\,dt$ $\leq C\|\sqrt{\rho_{0}}\theta_{0}\|_{L^{2}}^{2}+C\sup_{t\in[0,T]}\big(\|\nabla u(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}}\big)\int_{0}^{T}\|\sqrt{\rho}\dot{u}\|_{L^{2}}^{2}\,dt$
$\quad+C\sup_{t\in[0,T]}\big(\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}}\big)\int_{0}^{T}\|\nabla^{2}b\|_{L^{2}}^{2}\,dt$
$\quad+C\sup_{t\in[0,T]}\big[\|\nabla b(t)\|_{L^{2}}^{2}(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2})\big]\|\nabla b(t)\|_{L^{2}}^{2}.$ (3.15)
@@
$\quad+C\|b\|_{L^{2}}^{\frac{1}{2}}\|\nabla b\|_{L^{2}}^{\frac{1}{2}}\|\nabla^{2}b\|_{L^{2}}\|\nabla\theta\|_{L^{2}}+C\|\nabla\theta\|_{L^{2}}\|\nabla u\|_{L^{2}}\|\nabla b\|_{L^{2}}^{2}.$ (3.20)
@@
To control the last term of (3.22), Lemma 3.1 gives that
$\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\|\nabla b\|_{L^{2}}^{4}\,dt$ $\leq C\sup_{t\in[0,T]}\big(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2}\big)\|\nabla b(t)\|_{L^{2}}^{4}$
$\quad+C\sup_{t\in[0,T]}\big(\|\rho(t)\|_{L^{3}}^{2}\|\nabla b(t)\|_{L^{2}}^{4}\big)\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}\,dt$
$\leq C\sup_{t\in[0,T]}\big[\|\nabla b(t)\|_{L^{2}}^{2}(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2})\big]\|\nabla b(t)\|_{L^{2}}^{2}$
$\quad+C\sup_{t\in[0,T]}(\|\nabla b(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}})^{2}\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}\,dt.$
Choosing $\epsilon$ sufficiently small in (3.2) such that
$C\sup_{t\in[0,T]}(\|\nabla b(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}})^{2}\leq\frac{1}{2},$ (3.23)
integrating (3.22) over $(0,T)$, and using (3.23), we obtain the desired (3.3). ∎
@@
$\leq C\big(\|\nabla b_{0}\|_{L^{2}}^{2}+\|b_{0}\|_{L^{4}}^{4}\big)+C\sup_{t\in[0,T]}\big[(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})(\|\nabla u(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}})^{2}\big]\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}dt$
$\quad+C\sup_{t\in[0,T]}\big[(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2})\|\nabla u(t)\|_{L^{2}}^{2}\big]\|\nabla u(t)\|_{L^{2}}^{2}.$ (3.36)
@@
Combining (3) and (3.38), and using the a priori assumption (3.2) to choose $\epsilon$ sufficiently small such that
@@
Consequently, integrating (3.40) in $t$ over $(0,T)$ leads to
$\leq C\big(\|\nabla b_{0}\|_{L^{2}}^{2}+\|b_{0}\|_{L^{4}}^{4}\big)+C\sup_{t\in[0,T]}\big[(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})\|\nabla u(t)\|_{L^{2}}^{4}\big]\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}dt.$ (3.41)
For the last term in (3), it follows from (3.1) that
$C\sup_{t\in[0,T]}\big[\big(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2}\big)\|\nabla u(t)\|_{L^{2}}^{4}\big]\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}dt$
$\leq C\sup_{t\in[0,T]}\big[(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2})\|\nabla u(t)\|_{L^{2}}^{2}\big]\|\nabla u(t)\|_{L^{2}}^{2}$
$\quad+C\sup_{t\in[0,T]}\big[(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})(\|\nabla u(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}})^{2}\big]\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}dt,$
which together with (3) yields the desired (3.5). ∎
@@
$\quad+C\sup_{t\in[0,T]}\big(\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}}\big)\int_{0}^{T}\|\nabla^{2}b\|_{L^{2}}^{2}dt$
$\quad+C\sup_{t\in[0,T]}\big[\|\nabla b(t)\|_{L^{2}}^{2}(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2})\big]\|\nabla b(t)\|_{L^{2}}^{2}$
$\quad+C\sup_{t\in[0,T]}\big[(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2})\|\nabla u(t)\|_{L^{2}}^{2}\big]\|\nabla u(t)\|_{L^{2}}^{2}$
$\quad+C\sup_{t\in[0,T]}\big[(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})(\|\nabla u(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}})^{2}\big]\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}dt.$ (3.46)
@@
$C\max\big\{$ $\,\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}},\ \|\nabla b(t)\|_{L^{2}}^{2}\big(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2}\big),$
$\,\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2},\ \|\nabla u(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}}\big\}\leq\frac{1}{16},$ (3.47)
@@
$\|\sqrt{\rho}\theta(t)\|_{L^{2}}^{2}+\|\nabla b\|_{L^{2}}^{2}+\|b\|_{L^{4}}^{4}+\int_{0}^{T}\big(\|b_{t}\|_{L^{2}}^{2}+\|\nabla^{2}b\|_{L^{2}}^{2}+\||b||\nabla b|\|_{L^{2}}^{2}+\|\nabla\theta\|_{L^{2}}^{2}\big)dt$
@@
$\quad+C\sup_{t\in[0,T]}\big[(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2})\|\nabla u(t)\|_{L^{2}}^{2}\big]\|\nabla u(t)\|_{L^{2}}^{2}.$ (3.48)
@@
$\bar{\rho}\|\sqrt{\rho}\theta(t)\|_{L^{2}}^{2}+\bar{\rho}\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}\,dt$ $\leq C\bar{\rho}\|\sqrt{\rho_{0}}\theta_{0}\|_{L^{2}}^{2}+C\bar{\rho}\sup_{t\in[0,T]}\left(\|\nabla u(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}}\right)\int_{0}^{T}\|\sqrt{\rho}\dot{u}\|_{L^{2}}^{2}dt$
$\quad+C\bar{\rho}\sup_{t\in[0,T]}(\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}})\int_{0}^{T}\|\nabla^{2}b\|_{L^{2}}^{2}dt$
$\quad+C\bar{\rho}\sup_{t\in[0,T]}\big[\|\nabla b(t)\|_{L^{2}}^{2}(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2})\big]\|\nabla b(t)\|_{L^{2}}^{2}.$ (3.49)
@@
$\quad+C\bar{\rho}\sup_{t\in[0,T]}(\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}})\int_{0}^{T}\|\nabla^{2}b\|_{L^{2}}^{2}dt$
$\quad+C\bar{\rho}\sup_{t\in[0,T]}\big[\|\nabla b(t)\|_{L^{2}}^{2}(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2})\big]\|\nabla b(t)\|_{L^{2}}^{2}$
$\quad+C\sup_{t\in[0,T]}\big[(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2})\|\nabla u(t)\|_{L^{2}}^{2}\big]\|\nabla u(t)\|_{L^{2}}^{2}$
$\quad+C\sup_{t\in[0,T]}K_{1}(t)\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}dt+C\int_{0}^{T}K_{2}(t)dt$
$\leq C\left(\|\nabla u_{0}\|_{L^{2}}^{2}+(\bar{\rho}+1)\|\sqrt{\rho_{0}}\theta_{0}\|_{L^{2}}^{2}+\|\nabla b_{0}\|_{L^{2}}^{2}+\|b_{0}\|_{L^{4}}^{4}\right)+C\int_{0}^{T}K_{2}(t)dt,$ (3.50)
@@
$C\max\big\{$ $(\bar{\rho}+1)\|\nabla u(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}},~\bar{\rho}\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}},~\bar{\rho}\|\nabla b(t)\|_{L^{2}}^{2}(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2}),$
$\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2},~(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2})\|\nabla u(t)\|_{L^{2}}^{2}\big\}\leq\frac{1}{16}.$ (3.51)
In addition,
@@
$C\max\big\{$ $(\bar{\rho}+1)\|\nabla b(t)\|_{L^{2}}^{2}\big(\|\sqrt{\rho}u(t)\|_{L^{2}}^{2}+\|b(t)\|_{L^{2}}^{2}\big),~(\bar{\rho}+1)\|\nabla b(t)\|_{L^{2}}\|b(t)\|_{L^{2}},$
$\|\rho(t)\|_{L^{3}}\big(\|\nabla u(t)\|_{L^{2}}^{2}+\|\nabla b(t)\|_{L^{2}}^{2}\big)\big\}\leq\frac{1}{16}.$ (3.52)
@@
$\bigg(1+\bar{\rho}+\frac{1}{\bar{\rho}}\bigg)S_{t}^{1}S_{t}^{2}+(\bar{\rho}^{2}+\bar{\rho})\sup_{t\in[0,T]}\|\rho(t)\|_{L^{3}}\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}dt\leq C\bigg(1+\bar{\rho}+\frac{1}{\bar{\rho}}\bigg)S_{0}^{\prime}S_{0}^{\prime\prime}=C\epsilon_{0}\leq\frac{3}{2}\epsilon,$

### v2

@@ 1 Introduction
First, motivated by [13], we attempt to obtain the useful dissipation estimates from the basic energy inequality in terms of $L_{t}^{\infty}L_{x}^{3}$-norm of the density and $L_{t}^{2}L_{x}^{2}$-norm of $\nabla\theta$ (see (3.1)). With the help of a priori hypothesis (3.2) and the structure of system (1.1), we derive an estimate on the $L_{t}^{\infty}L_{x}^{3}$-norm of density (see Lemma 3.2). The next step is to obtain an estimate on the $L_{t}^{2}L_{x}^{2}$-norm of the gradient of temperature. At this stage, our argument relies on the temperature equation instead of the total energy equation used in [17]. This requires us to control the quantities $\sqrt{\rho}\dot{u}$ and $\nabla^{2}b$ in $L_{t}^{2}L_{x}^{2}$ (see Lemma 3.3). Multiplying (1.1)2 by $u_{t}$ and applying delicate energy estimates, the estimate on the $L_{t}^{2}L_{x}^{2}$-norm for $\sqrt{\rho}\dot{u}$ can be obtained. It should be emphasized that, due to the term $\operatorname{curl}b\times b$, the estimate for $\sqrt{\rho}\dot{u}$ generates the complicated quantity $K_{2}(t)$ (see (3.4)). To control all these magnetic effects, one needs appropriate estimates for the magnetic field, which are derived from (1.1)4 (see Lemma 3.5).
Next, the key issue is to enclose the bootstrap argument. Based on the estimates derived in Lemmas 3.13.5, one obtains the uniform bound of density (via characteristic arguments) and the desired control of the quantity $S_{t}$ provided that the initial scaling-invariant quantity $N_{0}$ is sufficiently small (see Corollaries 3.1 and 3.2). In particular, in order to construct a complete scaling-invariant quantity that meets the requirements of the a priori estimates and the bootstrap argument, we multiply (3) by $(1+\bar{\rho}+\tfrac{1}{\bar{\rho}})$. Moreover, we succeed in establishing the estimates for $\|\rho\|_{L^{3}}$ in (3) and $\|\nabla u\|_{L^{2}}^{2}$ in (3). During this process, because of the strong nonlinear coupling between the fluid velocity, temperature, and magnetic field, the key techniques in [23] cannot be applied in a straightforward manner. A new ingredient in our analysis is the introduction of the quantity
$(\bar{\rho}^{2}+\bar{\rho})\sup_{t\in[0,T]}\|\rho(t)\|_{L^{3}}\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}dt+(\bar{\rho}+1)\sup_{t\in[0,T]}\|\nabla b(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}dt+\sup_{t\in[0,T]}\|\nabla u(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt,$
which is assumed to be sufficiently small (see (3.2)). This structure allows us to completes the bootstrap argument when the condition (1.8) holds.
@@ 4. $(d)$
Under the hypotheses of Theorem 1.1, $\epsilon_{0}>0$ is the small positive constant that appears in (1.8). Then there exist positive constants $C_{1}$ and $\epsilon_{2}$, depending only on the physical parameters (but independent of the initial data, $\bar{\rho}$, and $T$), with $\epsilon_{2}$ further being determined by (3.14), (3.21), (3.33), (3.42), (3.45), (3), and (3.50), such that, if for all $t\in[0,T]$,
$\begin{cases}\rho\leq 2\bar{\rho},\\[5.69054pt] S_{t}+(\bar{\rho}^{2}+\bar{\rho})\sup_{t\in[0,T]}\|\rho(t)\|_{L^{3}}\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}\,dt+(\bar{\rho}+1)\sup_{t\in[0,T]}\|\nabla b(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt\\[5.69054pt] +\sup_{t\in[0,T]}\|\nabla u(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt\leq 2\epsilon,\end{cases}$ (3.2)
@@
$\begin{cases}\rho\leq\frac{3}{2}\bar{\rho},\\[5.69054pt] S_{t}+(\bar{\rho}^{2}+\bar{\rho})\sup_{t\in[0,T]}\|\rho(t)\|_{L^{3}}\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}\,dt+(\bar{\rho}+1)\sup_{t\in[0,T]}\|\nabla b(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt\\[5.69054pt] +\sup_{t\in[0,T]}\|\nabla u(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt\leq\frac{3}{2}\epsilon.\end{cases}$ (3.3)
@@
$\|\sqrt{\rho}\theta(t)\|_{L^{2}}^{2}+\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}\,dt$
$\leq C\|\sqrt{\rho_{0}}\theta_{0}\|_{L^{2}}^{2}+C\sup_{t\in[0,T]}\big(\|\nabla u(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}}\big)\int_{0}^{T}\|\sqrt{\rho}\dot{u}\|_{L^{2}}^{2}\,dt$
$\quad+C\sup_{t\in[0,T]}\big(\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}}\big)\int_{0}^{T}\|\nabla^{2}b\|_{L^{2}}^{2}\,dt+C\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\|\nabla b\|_{L^{2}}^{4}\,dt.$ (3.15)
@@
$\quad+C\|b\|_{L^{2}}^{\frac{1}{2}}\|\nabla b\|_{L^{2}}^{\frac{1}{2}}\|\nabla^{2}b\|_{L^{2}}\|\nabla\theta\|_{L^{2}}+C\|\nabla\theta\|_{L^{2}}\|\nabla u\|_{L^{2}}\|\nabla^{2}b\|_{L^{2}}^{2}.$ (3.20)
@@
Integrating (3.22) in $t$ over $(0,T)$, we obtain the desired (3.3). ∎
@@
$\leq C\big(\|\nabla b_{0}\|_{L^{2}}^{2}+\|b_{0}\|_{L^{4}}^{4}\big)+C\sup_{t\in[0,T]}(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})\sup_{t\in[0,T]}\|\nabla u(t)\|_{L^{2}}^{4}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt.$ (3.35)
@@
Combining (3) and (3.37), and using the a priori assumption (3.2) to choose $\epsilon$ sufficiently small such that for any $t$
@@
Consequently, integrating (3.39) in $t$ over $(0,T)$ leads to (3.5). ∎
@@
$\quad+C\sup_{t\in[0,T]}\big(\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}}\big)\int_{0}^{T}\|\nabla^{2}b\|_{L^{2}}^{2}dt+C\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\|\nabla b\|_{L^{2}}^{4}\,dt$
$\quad+C\sup_{t\in[0,T]}(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})\sup_{t\in[0,T]}\|\nabla u(t)\|_{L^{2}}^{4}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt.$ (3.44)
@@
$C\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}}\leq\frac{1}{16},$ (3.45)
@@
$\quad+C\sup_{t\in[0,T]}(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})\sup_{t\in[0,T]}\|\nabla u(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt\sup_{t\in[0,T]}\|\nabla u(t)\|_{L^{2}}^{2}$
$\quad+C\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\|\nabla b\|_{L^{2}}^{4}\,dt.$ (3.46)
@@
$\bar{\rho}\|\sqrt{\rho}\theta(t)\|_{L^{2}}^{2}+\bar{\rho}\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}\,dt$
$\leq C\bar{\rho}\|\sqrt{\rho_{0}}\theta_{0}\|_{L^{2}}^{2}+C\bar{\rho}\sup_{t\in[0,T]}\left(\|\nabla u(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}}\right)\int_{0}^{T}\|\sqrt{\rho}\dot{u}\|_{L^{2}}^{2}\,dt$
$\quad+C\bar{\rho}\sup_{t\in[0,T]}\big(\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}}\big)\int_{0}^{T}\|\nabla^{2}b\|_{L^{2}}^{2}\,dt+C\bar{\rho}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\|\nabla b\|_{L^{2}}^{4}\,dt.$ (3.47)
@@
$\quad+C\sup_{t\in[0,T]}(\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2})\sup_{t\in[0,T]}\|\nabla u(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt\sup_{t\in[0,T]}\|\nabla u(t)\|_{L^{2}}^{2}$
$\quad+C(\bar{\rho}+1)\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\|\nabla b\|_{L^{2}}^{4}\,dt+C\sup_{t\in[0,T]}K_{1}(t)\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}dt$
$\quad+C\bar{\rho}\sup_{t\in[0,T]}\big(\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}}\big)\int_{0}^{T}\|\nabla^{2}b\|_{L^{2}}^{2}\,dt+C\int_{0}^{T}K_{2}(t)dt$
$\leq C\left(\|\nabla u_{0}\|_{L^{2}}^{2}+(\bar{\rho}+1)\|\sqrt{\rho_{0}}\theta_{0}\|_{L^{2}}^{2}+\|\nabla b_{0}\|_{L^{2}}^{2}+\|b_{0}\|_{L^{4}}^{4}\right)+C\int_{0}^{T}K_{2}(t)dt$
$\quad+C(\bar{\rho}+1)\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\|\nabla b\|_{L^{2}}^{4}\,dt,$ (3.48)
@@
$C\max\big\{$ $(\bar{\rho}+1)\|\nabla u(t)\|_{L^{2}}^{2}\|\rho(t)\|_{L^{3}},~\bar{\rho}\|b(t)\|_{L^{2}}\|\nabla b(t)\|_{L^{2}},~\|b(t)\|_{L^{2}}^{2}\|\nabla u(t)\|_{L^{2}}^{2},$
$\sup_{t\in[0,T]}\|\nabla u(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt\big\}\leq\frac{1}{16}.$ (3.49)
In addition, noting that
$C(\bar{\rho}+1)\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\|\nabla b\|_{L^{2}}^{4}\,dt\leq C(\bar{\rho}+1)\sup_{t\in[0,T]}\|\nabla b(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt\sup_{t\in[0,T]}\|\nabla b(t)\|_{L^{2}}^{2},$
and
@@
$C\max\Big\{\|\rho(t)\|_{L^{3}}\|\nabla u(t)\|_{L^{2}}^{2},~(\bar{\rho}+1)\|\nabla b(t)\|_{L^{2}}\|b(t)\|_{L^{2}},~(\bar{\rho}+1)\sup_{t\in[0,T]}\|\nabla b(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt\Big\}\leq\frac{1}{16}.$ (3.50)
@@
$\bigg(1+\bar{\rho}+\frac{1}{\bar{\rho}}\bigg)S_{t}^{1}S_{t}^{2}+(\bar{\rho}^{2}+\bar{\rho})\sup_{t\in[0,T]}\|\rho(t)\|_{L^{3}}\int_{0}^{T}\|\nabla\theta\|_{L^{2}}^{2}dt$
$\quad+(\bar{\rho}+1)\sup_{t\in[0,T]}\|\nabla b(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt+\sup_{t\in[0,T]}\|\nabla u(t)\|_{L^{2}}^{2}\int_{0}^{T}\|\nabla u\|_{L^{2}}^{2}\,dt$
$\leq C\bigg(1+\bar{\rho}+\frac{1}{\bar{\rho}}\bigg)S_{0}^{\prime}S_{0}^{\prime\prime}=C\epsilon_{0}\leq\frac{3}{2}\epsilon,$